# True Zero-Shot Inference of Dynamical Systems Preserving Long-Term Statistics

### v1

@@ 4.3 Dynamical similarity
DynaMix’ mixture-of-experts design produces a natural similarity measure for comparing the dynamics of different systems, by evaluating the relative contribution of each of the experts to the forecast dynamics. As shown in Fig. 7a (bottom), different experts specialize in specific dynamical patterns or regimes. Hence, we can utilize the average expert usage $\bm{e}=\frac{1}{T}\sum_{t=1}^{T}\bm{w}^{exp}_{t}$ across the generated time series to construct a similarity metric
$S=\frac{1}{S_{max}}\left(||\bm{e}_{1}-\bm{e}_{2}||_{1}\right)^{-1}\;,$ (7)
defined here as the inverse $L_{1}$ norm of the $\bm{e}$ vectors between any two systems, rescaled to the range $[0,1]$.
Figure 7: a) Top: Vector fields of different experts which in combination determine the attractor. Bottom plot illustrates the dominant expert in each time window. b) Similarity matrix among all $84$ types of attractor dynamics in our dataset, determined by computing $S$ from the reconstructed time series. c) Examples of two pairs of attractors classified as similar (top), and one dissimilar pair.
@@ 4.4 Time series forecasting from a DS perspective
Figure 8: Time series forecasts for different data domains (gray: context, blue: ground truth, other colors: various model predictions): Forecasts of a) partially ($1$d) observed Lorenz-63 DS, b) electricity transformer temperature data, c) hourly car traffic data with weekly cycle, d) Huawei cloud request data, e) soil temperature development, f) air pressure data, g) human functional magnetic resonance imaging (fMRI) data, h) human electroencephalogram (EEG) data.
@@
System DynaMix Chronos-t5-base Mamba4Cast TTM
Partially obs. DS 0.11 0.25 0.32 0.11 0.32 0.39 8.14 0.35 0.76 6.54 0.46 0.89
ETTh1 0.8 0.08 5.97 0.33 0.09 3.22 0.94 0.1 3.63 6.15 0.09 4.64
Traffic 0.68 0.21 8.42 0.35 0.1 3.16 1.4 0.34 7.68 1.77 0.25 5.25
Cloud Requests 0.17 0.14 147.28 1.57 0.3 557.35 7.91 0.16 282.95 1.39 0.26 147.54
Weather (Temp.) 0.59 0.09 2.24 4.72 0.22 4.14 6.96 0.58 9.39 5.39 0.3 2.76
Weather (Press.) 0.41 0.19 8.85 6.62 0.23 9.05 7.03 0.52 8.6 7.27 0.25 10.03
Human fMRI 0.63 0.09 0.56 5.77 0.12 0.85 7.99 0.11 1.35 6.66 0.21 0.9
Human EEG 0.78 0.23 1.09 10.17 0.24 1.1 6.0 0.62 1.12 2.48 0.37 0.92
@@
We next tested and compared DynaMix on a variety of real-world datasets often used to probe time series models. As discussed in sect. 3.3, for DynaMix we used a multivariate embedding of the typically $1d$ empirical time series, to better reflect the underlying DS properties. Fig. 8a confirms for a $1d$ time series from the $3d$ chaotic Lorenz-63 system that this works well and DynaMix, in contrast to all other models, is still able to reconstruct the chaotic behavior. When applied to real-world time series, such as traffic, weather, or cloud request data, where previous TS foundation models have shown inconsistent results [80], our model produces accurate forecasts which reflect the dynamics well (Fig. 8b-f, see Appx. A.2 for further details on the datasets). Our model seems to particularly excel on human physiological signals, such as functional magnetic resonance imaging (fMRI) or electroencephalographic (EEG) recordings, where all other models essentially produce meaningless dynamics, at best capturing the mean trend (Fig. 8g,h).
@@
Table 1 quantitatively confirms that DynaMix mostly outperforms the other models tested on the DSR measures. Surprisingly, it often even outperforms them in terms of the forecasting error MAE (mean-absolute error), although its training corpus consisted purely of the simulated DS shown in Fig. 9, i.e. did not include any empirical data at all, let alone traffic, cloud usage, electricity, or temperature data! This is in stark contrast to the other TS foundation models tested here, which mostly had examples of such empirical data in their training corpus and hence might be expected to have an edge over DynaMix. Many real-world data, like weather, human fMRI, or EEG data, are known to bear signatures of deterministic chaos [56, 48], potentially at least partly explaining DynaMix’ strongly competitive forecasts, although it has never seen such type of data in training.
@@ 5 Conclusions
Here we introduce the first DSR foundation architecture that achieves zero-shot reconstructions of novel DS, accurately reproducing their long-term statistics, just from a provided context signal without any retraining or fine-tuning. It hugely outperforms current TS foundation models like Chronos in terms of DSR quality, although much more lightweight with only a fraction (0.1%) of the number of parameters. Consequently, DynaMix features inference times about an order of magnitude faster than the closest competitor. Most surprisingly, it even outperforms current TS foundation models on empirical data of types it has – in contrast to its competitors – never seen in training, often producing also better short-term forecasts. All of this is achieved with a fairly narrow training corpus consisting mainly of $3d$ synthetic data from chaotic systems, which suggests there may be considerable room for further improvement by extending DynaMix’ training to include, e.g., various empirical time series.
One important take-home from this work therefore is that foundation models built based on principles of DS theory may be able to profoundly improve performance of current zero-shot TS forecasters. Most (if not all) empirically observed time series come from some underlying DS, and acknowledging this fact in model training and construction may help to advance the field. Besides DynaMix’ specific DSR architecture, control-theoretically motivated training techniques, and embedding principles, also the fact that it has seen many chaotic systems in training may play a role. Most complex DS in nature [15, 82, 57], engineering [74], and society [54, 87] are likely chaotic, featuring an infinite spectrum of potential temporal patterns (unstable periodic orbits; [36]).
@@
For now, we have mainly focused on rather stationary synthetic and empirical time series (but see Fig. 8b). Changes in statistical properties over time, discontinuities, or widely differing time scales in the data observed, impose severe difficulties for zero-shot forecasting, as illustrated in Fig. 23a,b, which also plague other TS foundation models. However, this may partly be amended by explicitly including non-stationary DS or DS evolving on many different time scales, featuring sharp transitions, etc., in the training corpus. Incorrect embeddings, which may not be immediately obvious, can also lead to forecast failures (Fig. 23c).
@@ A.1 Methodological details
To this we add a regularization term meant to encourage the model to explore a wider region of state space based on the context, by enhancing the variance $\bm{\Sigma}$ of the exploration noise in the state attention:
$\mathcal{L}_{reg}=\lambda\left(1+\sum_{i}\Sigma_{ii}\right)^{-1}\;,$ (10)
where we chose $\lambda=0.01$. Rectified adaptive moment estimation (RADAM) [50] was employed as the optimizer, with $L=50$ batches of $S_{B}=16$ sequences per epoch, each of length $T=550$, and $2000$ epochs in total. We used a learning rate exponentially decaying from $\eta_{\text{start}}=2\cdot 10^{-3}$ to $\eta_{\text{end}}=10^{-5}$. The context length used in training was set to $T_{C}=500$, and the window of overlap with the model-generated time series to $\Delta t=50$, see sect. 3.2. Training was performed on a single CPU (18-Core Xeon Gold 6254), and a single epoch took $60$ seconds depending on sequence length and model size. At test time, no retraining or fine-tuning is performed, but utilizing the context $\bm{C}$ the model just forward-iterates from the last context time step. Hyperparameters for training were partly selected according to previous results in [7, 6], and partly a few different parameter settings were tested, namely $\lambda=\{0.0,0.01,0.02\}$, $\Delta t=\{0,50,100\}$, and $T_{C}=\{250,500\}$, as extensive grid search was not necessary to obtain a well-performing model.
@@
We use $J=20$ AL-RNN experts for our model. Each expert has a latent dimension of $M=20$, of which $P=7$ are rectified-linear units (ReLUs). As suggested in [7], we only use the first $N$ of the $M-P$ linear units for the readout. We followed the initialization protocol in [6, 77], drawing $\bm{W}$ from a Gaussian with mean $\bm{0}$ and $\bm{\sigma}=0.01\mathds{1}$, setting $\bm{h}=\mathbf{0}$, and $\bm{A}$ to the diagonal of a normalized positive-definite random matrix. The initial latent state was estimated as
@@
[…]
@@
We set $m=30$ bins per dimension, following [37]. To ensure a steady-state distribution is reached, long trajectories of $T=10,000$ time steps are sampled.
@@
[…]
@@
The $n$-step ahead prediction error to assess short-term forecast quality (with $n=10$ in our evaluation) is given by
$\text{PE}(n)=\left\|\bm{x}_{T_{C}+n}-\hat{\bm{x}}_{T_{C}+n}\right\|_{1}.$ (15)
The MAE for evaluating short term predictions on the empirical time series [39] is given by
@@
where we chose $n$ according to the data’s temporal scale and resolution ({ETTh1,fMRI}: $n=40$,{traffic, cloud requests, partially obs. DS}: $n=80$, {human EEG, weather air pressure}: $n=120$, {weather temperature}: $n=200$).
@@
For numerically estimating the maximum Lyapunov exponent $\lambda_{\text{max}}$ from a one-dimensional empirical time series $\{x_{t}\},t=1\dots T$, one common approach is the Rosenstein algorithm [68]. First, the time series is embedded into a state space using time-delay embedding as in eq. 5 (here using the PECUZAL algorithm [47]). For each embedded point $\bm{x}_{i}$, the nearest neighbor $\bm{x}_{j}$ is found under the constraints $|i-j|>l_{t}$, with $l_{t}$ a threshold to remove purely temporal neighbors (living on the same piece of trajectory), and a minimal initial state space separation $\|\bm{x}_{i}-\bm{x}_{j}\|_{2}>l_{s}$, where we chose $l_{t}=150$ (according to the mean periodicity [68]) and $l_{s}=0.2$. The local divergence between initially nearby trajectories is then tracked over a range of time steps $k=0\dots k_{max}$,
@@
The maximum Lyapunov exponent is given by
@@ A.2 Datasets
DynaMix is trained on about $1.6$ million simulated time series of length $T=550$ sampled from $34$ different $3d$ DS with cyclic or chaotic attractors, collected in [29]. Gaussian noise of $5\%$ of the data standard deviation was added to all dimensions, and time series were standardized to ensure comparable scaling of all data. Attractor dynamics were chosen to reflect different types of behavior, see Fig. 9.
Figure 9: Training data (blue) and their reconstruction in red.
@@
Our test set for DS consists of simulated time series of length $10^{5}$ sampled from $54$ different $3d$ DS collected in [29], which are not part of the training set. Furthermore, we evaluated the Selkov 2d DS [73] describing a kinetic model of an open monosubstrate enzyme reaction given by the equations
@@
where we chose $a=0.1$ and $b=0.5$.
For evaluation of time series forecasting, different real-world data were used. The traffic data are hourly recordings of the number of cars passing road junctions (https://www.kaggle.com/datasets/fedesoriano/traffic-prediction-dataset/data). The cloud data, also used in [80] to evaluate TS foundation models, are publicly available from Huawei Cloud https://github.com/sir-lab/data-release. It consists of function requests from Huawei’s serverless cloud platform. The weather data consists of daily sampled soil temperature and air pressure measured in the city of Mannheim, Germany (source: Deutscher Wetterdienst [German weather service]), and can be accessed via https://www.dwd.de/EN/ourservices/cdc/cdc_ueberblick-klimadaten_en.html. The functional magnetic resonance imaging (fMRI) data comes from human subjects performing cognitive tasks and is publicly available on GitHub [48]. We followed Kramer et al. [48] and selected the first principal component of BOLD activity in each of the $20$ brain regions. Kramer et al. [48] report a positive maximum Lyapunov exponent for models reconstructed from these time series, indicating their chaotic nature (see also [84]). The ETTh1 dataset is part of the Electricity Transformer Temperature (ETT) benchmark, which is widely used for evaluating TS forecasting models. It contains hourly data collected from a power transformer station [94] and can be accessed at https://github.com/zhouhaoyi/ETDataset. Electroencephalogram (EEG) data were taken from a study by Schalk et al. [71], comprising 64-channel data collected from human subjects performing various motor and imagery tasks. Following the approach of Brenner et al. [6], the signals were smoothed using a Hann window of length $15$. As for the training data, all test data were standardized for processing by the models.

### v2

@@ 4.3 Dynamical similarity
DynaMix’ mixture-of-experts design produces a natural similarity measure for comparing the dynamics of different systems, by evaluating the relative contribution of each of the experts to the forecast dynamics. As shown in Fig. 7a (bottom), different expert mixtures specialize in specific dynamical patterns or regimes. Hence, we can utilize the expert usage $\bm{w}^{exp}$ across the generated time series to construct a similarity metric $S=\frac{1}{1+W_{2}^{2}}\in[0,1]$, based on the squared 2-Wasserstein distance $W_{2}^{2}$ between the standardized expert weights $\bm{w}^{exp}$ of any two systems.
Figure 7: a) Top: Vector fields of different experts which in combination determine the attractor. Bottom plot illustrates the expert usage across time. b) Similarity matrix based on $S$ among all $84$ DS in the dataset, sorted via hierarchical clustering. Structure in the matrix indicates there are clusters of DS with high dynamical similarity. c) Examples of two pairs of attractors classified as similar (top), and one dissimilar pair (bottom).
@@ 4.4 Time series forecasting from a DS perspective
Figure 8: Comparison of DynaMix (red) to strongest competitor in terms of $D_{stsp}$ (other colors) on zero-shot forecasts of various empirical time series (see Fig. 27-30 for comparisons to all other TS foundation models): Forecasts of a) partially (11d) observed Lorenz-63 DS, b) electricity transformer temperature, c) hourly car traffic data with weekly cycle, d) Huawei cloud requests, e) soil temperature development, f) air pressure, g) human functional magnetic resonance imaging (fMRI), h) human electroencephalogram (EEG).
@@
System DynaMix Chronos-t5-base Mamba4Cast TTM TimesFM
$D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE
Partially obs. DS 0.02 0.25 0.39 0.02 0.32 0.39 8.35 0.35 0.76 4.76 0.31 0.90 7.30 0.36 0.71
ETTh1 0.22 0.08 4.92 0.23 0.09 3.22 0.80 0.10 3.63 6.36 0.09 4.64 0.24 0.10 3.04
Traffic 0.81 0.21 6.93 0.40 0.10 3.16 1.59 0.34 7.68 1.36 0.25 5.25 0.43 0.11 3.66
Cloud Requests 0.27 0.14 161.58 1.86 0.31 557.35 5.89 0.16 282.95 1.31 0.25 147.54 0.33 0.29 239.31
Weather (Temp.) 0.66 0.09 2.24 4.78 0.21 4.14 6.86 0.57 9.39 1.28 0.12 1.92 6.86 0.13 3.31
Weather (Press.) 0.39 0.19 9.03 6.68 0.23 9.05 7.07 0.52 8.60 7.31 0.26 10.03 9.59 0.60 8.41
Human fMRI 0.17 0.09 0.45 5.78 0.12 0.85 7.97 0.11 1.35 4.87 0.39 1.54 6.19 0.21 1.11
Human EEG 0.79 0.23 1.07 9.45 0.24 1.10 5.87 0.61 1.12 10.36 0.26 1.25 10.07 0.27 1.32
@@
We next tested and compared DynaMix on a variety of real-world datasets often used to probe time series models. As discussed in Sect. 3.3, for DynaMix we used an embedding of the typically $1d$ empirical time series, to better reflect the underlying DS properties (positional embedding for all empirical time series, akin to what is done in TS foundation models, and delay embedding for the Lorenz-63). Fig. 8a confirms for a $1d$ time series from the $3d$ chaotic Lorenz-63 system that this works well and DynaMix, in contrast to all other models, is still able to reconstruct the chaotic behavior. When applied to real-world time series, such as traffic, weather, or cloud request data, where previous TS foundation models have shown inconsistent results [84], our model produces accurate forecasts which reflect the dynamics well (Fig. 8b-f, see Appx. A.2 for further details on the datasets). Our model seems to excel particularly on human physiological signals, such as functional magnetic resonance imaging (fMRI) or electroencephalographic (EEG) recordings, where all other models essentially produce meaningless dynamics, at best capturing the mean trend (Fig. 8g,h).
@@
Table 1 quantitatively confirms that DynaMix mostly outperforms the other models tested on the DSR measures. Surprisingly, it often even outperforms them in terms of the forecasting error MAE, although its training corpus consisted purely of the simulated DS shown in Fig. 10, i.e. did not include any empirical data at all, let alone traffic, cloud usage, electricity, or temperature data! This is in stark contrast to the other TS foundation models tested here, which mostly had examples of such empirical data in their training corpus and hence might be expected to have an edge over DynaMix. Many real-world data, like weather, human fMRI, or EEG data, are known to bear signatures of deterministic chaos [58, 50], potentially at least partly explaining DynaMix’ strongly competitive forecasts, although it has never seen such type of data in training.
4.5 Which ingredients are most crucial for DynaMix’ performance? Ablation studies
Which are the components of the DynaMix framework contributing most to its success in zero-shot out-of-domain generalization? Training by STF (Fig. 26), a sufficiently diverse training corpus, a sufficient number of experts (Fig. 25), the attention mechanism, and context preprocessing by the CNN, all appear vital to its strong performance (Fig. 9). Replacing the final gating step (the MLP) by a simple linear operation and the exact expert model within the class of piecewise linear RNNs, however, had less of an impact. Similar results were obtained when the AL-RNN was swapped for the clipped shallow PLRNN [39], and a probabilistic version of the AL-RNN (adding a noise term $\bm{z}_{t}=F(\bm{z}_{t-1})+\bm{\eta},\bm{\eta}\sim\mathcal{N}(0,\Gamma)$) even slightly improved long-term performance, albeit at the cost of some short-term prediction accuracy. Using as experts LSTMs [41], vanilla RNNs or reservoir computers [65] instead led to decay in performance, in line with previous results establishing PLRNNs as SOTA DSR models which allow for efficient training by STF [6, 39, 7]. Hence, it appears that the whole ‘DSR package’ consisting of 1) architectural choices, 2) a training algorithm specialized for DSR, and 3) properties of the DS training corpus, is crucial to DynaMix’ success.
Figure 9: Ablation studies. a) Effect of removing or altering components of the gating network (no MLP, no CNN, no Attention, CNN to linear layer), including too few experts ($J<5$), replacing STF by standard BPTT training, restricting the training corpus to Lorenz-type systems only. b) Effect of replacing the AL-RNN by other RNN models. Error bars = STD
@@ 5 Conclusions
Here we introduce the first DSR foundation model that achieves zero-shot reconstructions of novel DS, accurately reproducing their long-term statistics, just from a provided context signal without any retraining or fine-tuning. It hugely outperforms current TS foundation models like Chronos in terms of DSR quality, although much more lightweight with only a fraction (0.1%) of the number of parameters. Consequently, DynaMix features inference times about an order of magnitude faster than the closest competitor. Most surprisingly, it even outperforms current TS foundation models on empirical data of types it has – in contrast to its competitors – never seen in training, often producing also better short-term forecasts. All of this is achieved with a fairly narrow training corpus consisting mainly of $3d$ synthetic data from chaotic systems, which suggests there may be considerable room for further improvement by extending DynaMix’ training to include, e.g., various empirical time series.
One important take-home from this work therefore is that foundation models built based on principles of DS theory may be able to profoundly improve performance of current zero-shot TS forecasters. Most (if not all) empirically observed time series come from some underlying DS, and acknowledging this fact in model training and construction may help to advance the field. Besides DynaMix’ specific DSR architecture, control-theoretically motivated training techniques, and embedding principles, also the fact that it has seen many chaotic systems in training may play a role. Most complex DS in nature [15, 86, 59], engineering [78], and society [56, 92] are likely chaotic. Moreover, many chaotic attractors, like the Lorenz-63, feature a skeleton of infinitely many (unstable) periodic orbits of all possible periods [36], thus inherently expressing a wide spectrum of temporal patterns.
@@
For now, we have mainly focused on rather stationary synthetic and empirical time series (but see Fig. 8b). Changes in statistical properties over time, tipping points, or widely differing time scales in the data, impose severe difficulties for zero-shot forecasting, as illustrated in Fig. 31a,b, which also plague other TS foundation models. This may partly be amended by explicitly including such types of non-stationary and multiscale DS in the training corpus. Another potential solution may be adding explicit filtering and decomposition modules, similar as in Auto- [95] or FEDformers [101]. A proof of concept of this idea is provided in Appx. A.5. Incorrect embeddings, which may not be immediately obvious, can also lead to forecast failures (Fig. 31c). Finally, continuous-time versions of the experts which can deal with sampling at irregular time points would be another fruitful direction.
@@ A.1 Methodological details
To this we add a regularization term meant to encourage the model to explore a wider region of state space based on the context, by enhancing the variance $\bm{\Sigma}$ of exploration noise in the state attention:
$\mathcal{L}_{reg}=\lambda\frac{1}{N}\sum_{i}\exp{\left(-\left|\Sigma_{ii}\right|/c\right)}\;,$ (9)
where we chose $\lambda=0.1$ and $c=0.01$. Rectified adaptive moment estimation (RADAM) [52] was employed as the optimizer, with $L=50$ batches of $S_{B}=16$ sequences per epoch, each of length $T=550$, and $2000$ epochs in total. We used a learning rate exponentially decaying from $\eta_{\text{start}}=5\cdot 10^{-3}$ to $\eta_{\text{end}}=10^{-5}$. The context length used in training was set to $T_{C}=500$, and the window of overlap with the model-generated time series to $\Delta t=50$, see Sect. 3.2. Training was performed on a single CPU (18-Core Xeon Gold 6254), and a single epoch took $30$ seconds depending on sequence length and model size. At test time, no retraining or fine-tuning is performed, but utilizing the context $\bm{C}$ the model just forward-iterates from the last context time step. Hyperparameters for training were partly selected according to previous results in [7, 6], and partly a few different parameter settings were tested, namely $\lambda=\{0.0,0.1,0.2\}$, $\Delta t=\{0,50,80\}$, and $T_{C}=\{250,500\}$, as extensive grid search was not necessary to obtain a well-performing model.
@@
We use $J=10$ AL-RNN experts for our model. Each expert has a latent dimension of $M=30$, of which $P=2$ are rectified-linear units (ReLUs). As suggested in [7], we only use the first $N$ of the MPM-P linear units for the readout. We followed the initialization protocol in [6, 81], drawing $\bm{W}$ from a Gaussian with mean $\bm{0}$ and $\bm{\sigma}=0.01\mathds{1}$, setting $\bm{h}=\mathbf{0}$, and $\bm{A}$ to the diagonal of a normalized positive-definite random matrix. The initial latent state was estimated as
@@
The gating network is implemented using a single-layer CNN with three channels and a kernel size of 2, stride of 1, and zero padding, with the identity as activation function. The MLP, which takes as inputs the concatenated weighted CNN outputs and current latent state, consists of two layers with ReLU activation. We initialize the temperature weights $\tau_{\text{att}}$ and $\tau_{\text{exp}}$ to $0.1$, the covariance matrix $\bm{\Sigma}=0.05\cdot\mathds{1}$, matrix $\bm{D}=\mathcal{I}_{(N\times M)}$, and draw all other CNN and MLP matrices from a Gaussian with standard deviation $0.01$. Model hyperparameters were selected according to previous results in [7] and by probing a few relevant parameter settings, namely $J=\{1,3,5,10,20,30\}$ (Fig. 25), $\text{CNN-layers}=\{1,2\}$, $\text{CNN-kernel-size}=\{2,3,5\}$, $\text{MLP-layers}=\{1,2,3\}$, and different activation functions $\{\mathds{1},\tanh{},\text{ReLU}\}$ for the MLP/CNN. We found that an extensive grid search was not necessary to obtain a well-performing model.
@@
We set $m=30$ bins per dimension, following [38]. For the $6d$ Lorenz-96 we use $m=5$ and for the empirical time series $m=20$. To ensure a steady-state distribution is reached, long trajectories of $T=10,000$ time steps are sampled from the DS.
@@
$H(F(\omega),G(\omega))=\sqrt{1-\int_{-\infty}^{\infty}\sqrt{F(\omega)G(\omega)}d\omega}\;\in[0,1],$ (13)
@@
For assessing short-term forecast quality aggregated across the whole test set of $54$ DS, we used a normalized $n$-step ahead prediction error (with $n=10$ in our evaluation) as recommended in [40], given by
$\text{MASE}=\frac{1}{N}\sum_{i=1}^{N}\frac{\dfrac{1}{n}\sum_{t=T_{C}+1}^{T_{C}+n}\left|x_{i,t}-\hat{x}_{i,t}\right|}{\dfrac{1}{T}\sum_{t=T_{C}+1}^{T_{C}+T}\left|x_{i,t}-x_{i,t-1}\right|}$ (14)
For individual comparisons on single empirical times, the MAE for evaluating short term predictions [40] was used:
@@
where we chose $n$ according to the data’s temporal scale and resolution ({ETTh1,fMRI}: n=40n=40,{traffic, cloud requests, partially obs. DS}: $n=80$, {human EEG, weather air pressure}: $n=120$, {weather temperature}: $n=200$).
@@
For numerically estimating the maximum Lyapunov exponent $\lambda_{\text{max}}$ from a one-dimensional empirical time series $\{x_{t}\},t=1\dots T$, one common approach is the Rosenstein algorithm [72]. First, the time series is embedded into a state space using time-delay embedding as in eq. 5 (e.g. using the PECUZAL algorithm [49]). For each embedded point $\bm{x}_{i}$, the nearest neighbor $\bm{x}_{j}$ is found under the constraints $|i-j|>l_{t}$, with $l_{t}$ a threshold to remove purely temporal neighbors (living on the same piece of trajectory), and a minimal initial state space separation $\|\bm{x}_{i}-\bm{x}_{j}\|_{2}>l_{s}$, where we chose $l_{t}=150$ (according to the mean periodicity [72]) and $l_{s}=0.25$. The local divergence between initially nearby trajectories is then tracked over a range of time steps $k=0\dots k_{max}$,
@@
An estimate of the maximum Lyapunov exponent is given by
@@ A.2 Datasets
DynaMix is trained on about $0.6$ million simulated time series of length $T=550$ sampled from $34$ different $3d$ DS with cyclic or chaotic attractors, collected in [29]. Time series were standardized to ensure comparable scaling of all data, and Gaussian noise of $5\%$ of the data standard deviation was added to all dimensions for the ground truth. Attractor dynamics were chosen to reflect different types of behavior, see Fig. 10.
Figure 10: Training data (gray) and their reconstructions in red.
@@
Our test set for DS consists of simulated time series of length $10^{5}$ sampled from $54$ different $3d$ DS collected in [29], which are not part of the training set.
Furthermore, we evaluated several $2d$ systems: The Selkov system [77] describing a kinetic model of an open monosubstrate enzyme reaction given by the equations
@@
where we chose $a=0.1$ and $b=0.5$. The Van-der-Pol system [87] describes self-sustaining oscillations in vacuum tubes by
$\frac{dx}{dt}$ $=y,$ (20)
$\frac{dy}{dt}$ $=\mu(1-x^{2})y-x,$
where we chose $\mu=0.5$. Similar to the Van-der-Pol system, the Rayleigh oscillator [61] describes a self-sustained nonlinear oscillator through
$\frac{dx}{dt}$ $=y,$ (21)
$\frac{dy}{dt}$ $=\mu(1-\frac{y^{2}}{3})y-x,$
where we chose $\mu=1.0$. As a higher dimensional DS test case we use the Lorenz-96 system [54] defined by
$\frac{\text{d}x_{i}}{\text{d}t}=(x_{i+1}-x_{i-2})x_{i-1}-x_{i}+F,$ (22)
with system variables $x_{i}$, i=1,,Ni=1,...,N, and forcing term $F$ (here, $F=8$ and $N=6$, in the chaotic regime). Furthermore, cyclic boundary conditions are assumed with $x_{-1}=x_{N-1}$, $x_{0}=x_{N}$, $x_{N+1}=x_{1}$, and the system was solved with integration step $\Delta t=0.08$.
We further probed forecasting and DSR on different types of real-world data: The traffic data are hourly recordings of the number of cars passing road junctions (https://www.kaggle.com/datasets/fedesoriano/traffic-prediction-dataset/data). The cloud data, also used in [84] to evaluate TS foundation models, are publicly available from Huawei Cloud (https://github.com/sir-lab/data-release). It consists of function requests from Huawei’s serverless cloud platform. The weather data consists of daily sampled soil temperature and air pressure measured in the city of Mannheim, Germany (source: Deutscher Wetterdienst [German weather service]), and can be accessed via https://www.dwd.de/EN/ourservices/cdc/cdc_ueberblick-klimadaten_en.html. The functional magnetic resonance imaging (fMRI) data comes from human subjects performing cognitive tasks and is publicly available on GitHub [50]. We followed Kramer et al. [50] and selected the first principal component of BOLD activity in each of the $20$ brain regions. Kramer et al. [50] report a positive maximum Lyapunov exponent for models reconstructed from these time series, indicating their chaotic nature (see also [89]). The ETTh1 dataset is part of the Electricity Transformer Temperature (ETT) benchmark, which is widely used for evaluating TS forecasting models. It contains hourly data collected from a power transformer station [100] and can be accessed at https://github.com/zhouhaoyi/ETDataset. Electroencephalogram (EEG) data were taken from a study by Schalk et al. [75], comprising 64-channel data collected from human subjects performing various motor and imagery tasks. Following the approach of Brenner et al. [6], the signals were smoothed using a Hann window of length $15$. As for the training data, all test data were standardized for processing by the models.